# Early Timestep Zero-Shot Candidate Selection for Instruction-Guided Image Editing

### v1

@@ 5.4 Ablation Study
Figure 10: Extracted masks at different timesteps. The right three columns show masks extracted at individual denoising steps ($t=95,90$, and $80$) for IP2P [1]. The leftmost column of masks shows the averaged mask over $t\in[80,100]$, which consistently yields more reliable results across diverse cases.
@@ Acknowledgments
This work was partly supported by multiple grants from the Institute of Information & Communications Technology Planning & Evaluation (IITP), funded by the Korean government (MSIT): [Grant No. RS2021-II211343, Artificial Intelligence Graduate School Program, Seoul National University] and [Grant No. RS-2024-00457882, AI Research Hub Project]. It was also supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) [Grant No. RS-2023-00209060, A Study on Optimization and Network Interpretation Method for Large-Scale Machine Learning]. Additionally, this research was conducted as part of the ”AI Media and Cultural Enjoyment Expansion” Project, supported by the Ministry of Science and ICT and the National IT Industry Promotion Agency (NIPA) in 2025.
@@ Appendix A Comparison with Existing Work
We provide a table comparing our work with previous image editing studies in Table 2. Our method is the first to introduce optimal seed selection for instruction-guided editing and uniquely enables MLLM-based instruction prompt selection, which is absent in existing approaches. Unlike prior methods, our ELECT framework achieves these capabilities without requiring external segmentation models or source/target prompt pairs.
@@
Table 2: Comparison of Methods Addressing Background Inconsistency in Text-guided Image Editing.
Appendix B Detailed Experimental Setup
Baselines. We establish 5 diffusion-based instruction-guided image editing models as baselines. All models operate under a constrained setting where they take only the source image and user instruction as inputs, without access to ground-truth masks or source/target prompts. The instruction-guided image editing models considered in this work include InstructPix2Pix [1], MagicBrush [50], InstructDiffusion [9], MGIE [7], and UltraEdit [53]. Among them, UltraEdit is a fine-tuned model based on Stable Diffusion 3, demonstrating that our method can also enhance the performance of Rectified-Flow models effectively.
@@ Appendix C Additional analysis
C.1 Analysis of Timestep for Selection
@@
(a) InstructPix2Pix (IP2P) [1]
(b) MagicBrush [50]
(c) InstructDiffusion (InsDiff) [9]
(d) MGIE [7]
(e) UltraEdit [53]
C.2 Analysis of Mask Extraction
@@ Appendix E ELECT for Instruction Prompt Selection
If either metric scores 0.0, the edit is classified as a failure, triggering the prompt selection process. Table 3 provides a comprehensive summary of the failure rates across different models and selection methods. The prompt used for the evaluation of the MLLM is as follows.
@@
Table 3: Failure case analysis using the MLLM[34] evaluator. We evaluated PIE-bench data based on Background Consistency (BC) and Instruction Following (IF), categorizing each as 0, 0.5, or 1.0. Total number of data is 700 in PIE-bench. A case was considered a failure if either score was 0. We set the number of seeds to $N=10$ for ELECT and applied prompt selection only to the remaining failed cases after seed selection, with $N=10$ prompts for re-selection. As a result, the editing failure rate significantly decreased, successfully correcting approximately 40% of previously failed baseline cases.
Model Seed Selection Method BC IF VIEScore (Semantic Consistency) ($\uparrow$)
@@
Prompt Selection via MLLM.
For failed cases, we introduce an additional step where an MLLM generates alternative instruction prompts. Given the input image and the original prompt, the MLLM is instructed to produce semantically equivalent but lexically varied instructions. To ensure diversity, we explicitly include constraints in the prompt, encouraging variations in wording, phrasing, and structure without altering the intended meaning (Algorithm 2).
This iterative process improves the likelihood of finding a prompt that falls within the model’s learned distribution, ultimately increasing the success rate of edits. The evaluation prompt and instruction generation prompt are provided below.
2:Best edited image $x^{*}$
3:$x^{0}\leftarrow\texttt{ELECT}(\mathbb{S},t_{\text{stop}})$triangleright Algorithm 1
4:if $\mathcal{M}_{\phi}(I,C_{T},x^{0},{\texttt{"evaluate $x^{0}$"}})>0$ then
5:     return $x^{*}\leftarrow x^{0}$triangleright Exit on edit success
6:end if
7:Sample a single initial noise $z_{T}\sim\mathcal{N}(0,I)$
8:$z_{T}^{1}=\cdots=z_{T}^{N}\leftarrow z_{T}$
9:$\{C_{i}\}_{i=1}^{N}\leftarrow\mathcal{M}_{\phi}(I,C_{T},\texttt{{"generate $N$prompts"}})$" )
11:     for $i\leftarrow 1,2,\dots,N$ do
12:         $z_{t-1}^{i}\leftarrow\texttt{Denoise}(z_{t}^{i},t,I,C_{i})$
13:     end for
14:end for
15:for $i\leftarrow 1,2,\dots,N$ do
17:end for
18:$i^{*}\leftarrow\arg\min_{i\in[N]}S^{\text{BIS}}(i,t_{\text{stop}})$triangleright Select best prompt
19:for $t=t_{\text{stop}}\to 1$ do $\triangleright$ Continue denoising $i^{*}$
20:     $z^{i^{*}}_{t-1}\leftarrow\texttt{Denoise}(z^{i^{*}}_{t},t,I,C_{i^{*}})$
21:end for

### v2

@@ 5.4 Ablation Study
Figure 8: Extracted masks at different timesteps. The right three columns show masks extracted at individual denoising steps ($t=95,90$, and 808080) for IP2P [1]. The leftmost column of masks shows the averaged mask over $t\in[80,100]$, which consistently yields more reliable results across diverse cases
@@ Acknowledgments
This work was supported by multiple grants from the Institute of Information & Communications Technology Planning & Evaluation (IITP), funded by the Korean government (MSIT): RS-2024-00457882 (AI Research Hub Project), RS-2025-25442338 (AI Star Fellowship Support Program, Seoul National University), RS-2025-02305581, and RS2021-II211343 (Artificial Intelligence Graduate School Program, Seoul National University). It was also supported by a grant from the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (RS-2025-02307233), and by the National Research Foundation of Korea (NRF) grants funded by MSIT: RS-2023-00209060 (A Study on Optimization and Network Interpretation Method for Large-Scale Machine Learning), RS-2023-00242443, and RS-2023-00282907. This research was also conducted as part of the ”AI Media and Cultural Enjoyment Expansion” Project, supported by the Ministry of Science and ICT and the National IT Industry Promotion Agency (NIPA) in 2025.
@@ Appendix A Comparison with Existing Work
We provide a table comparing our work with previous image editing studies in Tab. 2
Appendix B Detailed Experimental Setup
Baselines. We establish 5 diffusion-based instruction-guided image editing models as baselines. All models operate under a constrained setting where they take only the source image and user instruction as inputs, without access to ground-truth masks or source/target prompts. The instruction-guided image editing models considered in this work include InstructPix2Pix (IP2P) [1], MagicBrush [50], InstructDiffusion (InsDiff) [9], MGIE [7], and UltraEdit [53]. Among them, UltraEdit is a fine-tuned model based on Stable Diffusion 3, demonstrating that our method can also enhance the performance of Rectified-Flow models effectively.
(a) InstructPix2Pix
(b) MagicBrush
(c) InstructDiffusion
(d) MGIE
(e) UltraEdit
Table 2: Comparison of Methods Addressing Background Inconsistency in Text-guided Image Editing. Our method is the first to introduce optimal seed selection for instruction-guided editing and uniquely enables MLLM-based instruction prompt selection, which is absent in existing approaches. Unlike prior methods, our ELECT framework achieves these capabilities without requiring external segmentation models or source/target prompt pairs.
@@ Appendix C Additional Analysis
C.1 Motivation Validation
We ranked each sample’s seed outputs by background MSE—treating the background as the preserved region—and found that lower background MSE strongly correlates with higher edit quality and better instruction following. Fig. 10(a) quantifies this relationship, while a complementary user study in Fig. 10(b) further corroborates it. In that study, 34 participants evaluated 52 high-variance, category-balanced PIE-Bench samples, each presenting ten outputs (seeds 1–10) produced by IP2P, InsDiff, or UltraEdit. Participants tagged each image as well- or poorly-edited, and we defined user preference as the difference between positive and negative responses for each rank. Inter-rater agreement was substantial (Krippendorff’s $\alpha$=0.7535), underscoring the reliability of the findings. Moreover, Fig. 1 shows clear performance gains as background MSE decreases, confirming that background-consistent edits inherently yield superior results.
Figure 10: (a) Higher background consistency (lower BG-MSE rank) correlates with higher CLIPScore and VIEScore across models. (b) User Study: User preference strongly correlates with background consistency (Pearson $r=0.534$), confirming perceptual alignment.
C.2 Analysis of Timestep for Selection
@@
We further examine how the benefits of ELECT scale with the number of candidate seeds $N$. As shown in Fig. 11, the marginal performance improvements steadily taper off as $N$ grows, yet they remain consistently positive relative to the fixed-seed baseline. This figure extends the saturation trend observed in Fig. 6 to a broader range of $N$ values, confirming that larger candidate pools yield diminishing—but still meaningful—returns.
Figure 11: Performance trend in Fig. 6 when N=25, 50, 100.
C.3 Analysis of Mask Extraction
@@
Figure 12: ELECT performance variation with respect to stopping timestep ($t_{\text{stop}}$) with fixed number of seeds.
C.4 Analysis for Global Edits.
[…]
Figure 14: Qualitative results of global editing with ELECT.
C.5 Failure Cases.
Although ELECT can occasionally select edits that are overly mild—preserving too much of the background and dampening the intended change (see Fig. 15)—we observed that such instances were relatively uncommon in our experiments. This rarity likely stems from the strong modification bias of many instruction-guided image-editing models, which tends to push outputs toward more pronounced alterations, making excessive, unintended changes the more prevalent concern in practice.
Figure 15: A failure case of ELECT.
C.6 Another Signal for ELECT.
We also evaluated two alternative early-step cues for ELECT: foreground MSE (FG MSE) and CLIP-based text–image alignment. As illustrated in Fig. 16, both alternatives actually reduced performance—CLIP struggles with highly noisy early-timestep latents, and selecting the seed with the highest FG MSE often drives excessive foreground changes, yielding over-edited images. A simple hybrid rule mitigates these issues: we choose between the seed with the lowest BG MSE and the one with the highest FG MSE, whichever produces the better preliminary score. This strategy (yellow bars) enhances robustness, rescuing failure cases where pure background consistency alone falters (see Fig. 15). These observations highlight the opportunity to combine our pixel-wise background metric with complementary cues—such as foreground change or structural similarity—to support a more reliable, multi-aspect selection process. Realizing such a multi-objective framework will require deeper analysis, which we regard as a promising direction for future work.
Figure 16: Comparison with various signals for ELECT.
@@ Appendix E ELECT for Instruction Prompt Selection
2:Best edited image $x^{*}$
[…]
4:if $\mathcal{M}_{\phi}(I,C_{T},x^{0},{\texttt{"evaluate $x^{0}$"}})>0$ then
5:  return $x^{*}\leftarrow x^{0}$triangleright Exit on edit success
6:end if
7:Sample a single initial noise $z_{T}\sim\mathcal{N}(0,I)$
[…]
11:  for $i\leftarrow 1,2,\dots,N$ do
12:   $z_{t-1}^{i}\leftarrow\texttt{Denoise}(z_{t}^{i},t,I,C_{i})$
13:  end for
14:end for
15:for $i\leftarrow 1,2,\dots,N$ do
17:end for
[…]
19:for $t=t_{\text{stop}}\to 1$ do $\triangleright$ Continue denoising $i^{*}$
20:  $z^{i^{*}}_{t-1}\leftarrow\texttt{Denoise}(z^{i^{*}}_{t},t,I,C_{i^{*}})$
21:end for
@@
If either metric scores 0.0, the edit is classified as a failure, triggering the prompt selection process. We provide the useful prompt used for MLLM evaluation:
(keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help
@@
Prompt Selection via MLLM.
For failed cases, we introduce an additional step where an MLLM generates alternative instruction prompts (Fig. 17). Given the input image and the original prompt, the MLLM is instructed to produce semantically equivalent but lexically varied instructions. To ensure diversity, we explicitly include constraints in the prompt, encouraging variations in wording, phrasing, and structure without altering the intended meaning.
This iterative process improves the likelihood of finding a prompt that falls within the model’s learned distribution, ultimately increasing the success rate of edits. The instruction generation prompt are provided below:
(keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help
Quantitative Results.
We evaluated PIE-bench data based on Background Consistency (BC) and Instruction Following (IF), categorizing each as 0, 0.5, or 1.0. Total number of data is 700 in PIE-bench. A case was considered a failure if either score was 0. We set the number of seeds to $N=10$ for ELECT and applied prompt selection only to the remaining failed cases after seed selection, with $N=10$ prompts for re-selection. As a result, the editing failure rate significantly decreased, successfully correcting approximately 40% of previously failed baseline cases. (Tab. 3) Furthermore, we present the results of a comprehensive comparative evaluation of seed/prompt selection techniques across the whole metrics. (Tab. 4)
@@
Table 3: Failure case analysis using the MLLM[34] evaluator.
Model
Seed Selection
Method
BC IF VIEScore (Semantic Consistency) ($\uparrow$)